OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = 2x2 + 4x – 6?

OpenStudy (anonymous):

Put this equation in the standard form y = mx + b, where b will be your y intercept. To find your x intercept (which is the point whre the equation crosses the x-axis) you must set y = 0 because that's when the equation will cross the x-axis.

OpenStudy (anonymous):

Frkkk, that isn't a linear equation. If it's a quadratic you might want to use the quadratic formula.

OpenStudy (anonymous):

\[2x^2+4x-6=0\] \[x^2+2x-3=0\] \[(x+3)(x-1)=0\] \[x=-3, x=1\] are your x - intercepts y intercept is 6